OpenStudy (anonymous):

Prove that 2tan^-1 1/2 + tan^-1 1/7 = tan^-1 31/17

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

Sorry, I really don't know how to do this :/ hopefully @jdoe0001 @zepdrix @phi @shamim can help?

OpenStudy (anonymous):

I went through my book a bit, You gotta use tan^-1 x + tan^-1 y as tan^-1 x+y/1 + xy So its 2(1/2) + 1/7/1 + 1/2 x 1/7 1 + 1/7 / 1 + 1/14 Or 8/7/15/14 cancel cancel.. 16/15 tan^-1 16/15 Someone see if this is correct? :(

OpenStudy (anonymous):

@sleepyhead314 its okay <3

OpenStudy (sleepyhead314):

http://www.rapidtables.com/math/trigonometry/arctan.htm

OpenStudy (anonymous):

darn, its wrong. Its given in the question that its 2tan^-1 1/2 + tan^-1 1/7 So it means I gotta first get this in the form of tan^-1 x + tan^-1 y The 2tan^-1 x will be solved using 2x/1-x^2 that is 2 x 1/2 / 1 - (1/2)^2 = 4/3 so its 4/3 + 1/7 / 1 - 4/3 x 1/7 28 + 3/21 / 1 - 4/21 31/21 / 17/21 31/21 x 21/17 cancel cancel and 31/17 tan^-1 31/17 = tan^-1 31/17 Thus proved

OpenStudy (anonymous):

thank you! (: